One observer at point A estimates the angle of elevation to the basket of a hot air balloon to be 55°, while another observer 100 yards away at point B estimates the angle of elevation to be 36°. How high off the ground is the basket of the hot air balloon? Round to the nearest whole number.
Answer by TimothyLamb(4379)

---
make a triangle between basket and points A and B:
---
point A supplementary angle:
180 - 55 = 125
---
angle at point C, the basket:
180 - 125 - 36 = 19
---
use the law of sines to find the length of side "b":
---
b/sinB = c/sinC
b = sinB * c/sinC
b = sin(36) * 100/sin(19)
b = 180.541384114 yards
---
make a new right triangle with a vertical side "a" directly below the basket, point B as the 90 degree point, and point A as the other vertex:
---
the new angle at point C (the basket):
---
180 - 55 - 90 = 35
---
use the law of sines to find the length of side "a":
---
a/sinA = b/sinB
a = sinA * b/sinB
a = sin(55) * 180.541384114/sin(90)
a = 147.891 yards
---
answer:
height of basket above ground = 147.9 yards
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---